question: 65+n=85-16+3n


Process so far:65+n=69+3n
65+n-n=69+3n-n
65=69-2n
65-2n=69+2n-2n

Answer by checkley71(8403)

65+N=69+3N
N-3N=69-65
-2N=4
N=4/-2
N=-2 ANSWER.
PROOF:
65+(-2)=69+3(-2)
65-2=69-6
63=63